Write the equation in standard form:
y+2=1/2(x-3)
Thanks!

Found 2 solutions by OHtutor, stanbon:
Answer by OHtutor(16)

2(y+2)(x-3)=1
2(xy-3y+2x-6)=1
2xy-6y+4x-12=1
4x-6y+2xy=13

Answer by stanbon(48516)

y+2=1/2(x-3)
y+2=(1/2)x-(3/2)
y= (1/2)x-(3/2)-2
y=(1/2)x-(7/2)
Cheers,
Stan H.